OpenStudy (anonymous):

State the domain of the rational function. f(x)=(x-9)(x+2)/x^2-1

OpenStudy (anonymous):

all real numbers exccept -1 and 1 because x^2-1=0 x^2=1 sqrt x^2= plus minus sqrt 1 x= plus/minus 1

OpenStudy (anonymous):

Thanks so much! I was having difficulty with this one.